40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14653     Entry time: Tue Jun 4 10:56:31 2019     In reply to: 14647     Reply to this: 14659
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

I briefly managed to lock the IMC today - it stayed locked for ~10 minutes. Attachment #1 shows spectra of a few error and control signals for today's lock, and from a stretch yesterday before the problems surfaced*. The 60 Hz lines are much bigger, and MC_F signals broadband excess noise above a few Hz. I suspect a problem somewhere in the electronics.

*I confess the comparison isn't entirely valid because I had to tweak the FSS FAST gain from its nominal value of 22 to 25 in order to get the PC drive RMS down to the ~1.5V level. At the nominal gain setting, with the laser frequency locked to the cavity length, the PC Drive RMS was ~4 V. Still, indicative of something being off in the electronics.

Attachment 1: IMCdiag.pdf  61 kB  Uploaded Tue Jun 4 11:59:04 2019  | Hide | Hide all
IMCdiag.pdf
ELOG V3.1.3-